An item is regularly priced at $75 . It is now priced at a discount of 20% off the regular price. What is the price now?
Answer by ikleyn(52778)

.
The price now is 75 - 0.2*75 = 75 - 15 = 60 dollars.

-------------------

On percentage,  see the lessons
    - Percentage problems
    - Percentage word problems (Type 1 problems, Finding the Part)
    - Percentage word problems (Type 2 problems, Finding the Rate)
    - Percentage word problems (Type 3 problems, Finding the Base)
    - More complicated percentage problems
    - Problems on percentage that lead to unexpected results
in this site.

They are written specially for you.

Read them and become an expert in this area.